OpenStudy (anonymous):

3(4-t)^1/4=6^1/4. Please help to solve this step by step

OpenStudy (tkhunny):

#1 - Realize what cannot be! 4 - t >= 0. Why? #2 - Have you considered raising to the 4th power?

OpenStudy (anonymous):

That ^(1/4) is basically this : \[\sqrt[4]{x}\] If you take each side and put it to the 4th power then the radicals drop out and you are left with: 3(4-t)=6

OpenStudy (anonymous):

but the answer is 100/67. You got it wrong

OpenStudy (anonymous):

Really? Is it 3(4-t)^(1/4) or [3(4-t)]^(1/4)?

OpenStudy (anonymous):

it is 3 times 4th root of 4-t=4th root of 6